OpenStudy (anonymous):

What is the slope of the line that passes through (3, –7) and (–1, 1)? undefined 0 –2 2

OpenStudy (ash2326):

Slope of a line passing through two points (x1, y1) and (x2, y2) is \[\frac {y2-y1}{x2-x1}\] (x1, y1)=(3, -7) (x2, y2)=(-1, 1) Can you solve this @miah ?

OpenStudy (anonymous):

yeah,

OpenStudy (anonymous):

1.5 ?

OpenStudy (ash2326):

\(\frac{1-(-7)}{-1-3}\) \[\frac{8}{-4}\ne 1.5\]

OpenStudy (anonymous):

so, undefined ?

OpenStudy (ash2326):

\[\frac{-8}{4}=???\]

OpenStudy (ash2326):

I meant that you have did some wrong calculation

OpenStudy (anonymous):

-2

OpenStudy (ash2326):

Yes

OpenStudy (anonymous):

Thanks (: